OpenStudy (anonymous):

if log√(x^2+y^2) =tan^-1(y/x) prove that dy/dx=(x+y)/(x-y) m not getting teh answer :'(

hartnn (hartnn):

simplify the left side a lil bit \(\large \log \sqrt a= \log a^{1/2}=....?\)

OpenStudy (anonymous):